# Deep Self-Supervised Disturbance Mapping with the OPERA Sentinel-1 Radiometric Terrain Corrected SAR Backscatter Product

### v1

@@ VI-E Sensitivity to $\tau$
First, we evaluate the sensitivity of each method’s $F_{1}$ score to the threshold $\tau$ for the landslide event in Figure 17. Natively, the deep methods’ threshold is in units of standard deviations. For the log ratio, the threshold is in units of decibels. However, to compare all the methods side-by-side, we plot the thresholds on the x-axis as a percentage of the maximum threshold (where no pixels would be labeled positive). As described previously, the transformer achieves the best $F_{1}$ score. However, this plot illuminates that the transformer is also less sensitive to the threshold $\tau$ compared to the RNN and log ratio at their respective peaks.
Furthermore, in Figure 18, we show that the transformer achieves strong $F_{1}$ scores on all tasks at similar thresholds; a threshold of approximately $5$ standard deviations would result in $F_{1}$ scores around $0.6$ on all three tasks. While further validation is required, this is initial evidence that one could a priori select $\tau\approx 5$ and expect the transformer to produce high quality disturbance delineations across environments - even when ground truth is not available. This is a critical quality for a model to possess for operational deployment. Future work will focus on further corroborating this claim on more disturbance events.
@@ VI-F Ablation Studies
This subsection describes and empirically justifies some of the transformer and RNN hyperparameter choices made in this work.
@@ VI-F1 Input and Patch Size
Table II compares the performance of the same transformer-based model with different input and patch sizes. We report PR AUC. The architecture and training procedure for each model in the table is the same as reported in Section III-C.
@@
TABLE II: Comparison of the same transformer-based models with different input (I) and patch sizes (P). The patches are given a positional embedding in space and time. We report PR AUC for each disturbance mapping task. Best results are bolded. The patch size of 8 is clearly superior to 16, while the input size of 16 is marginally better than 32.
@@
TABLE III: Comparison of the same RNN-based model with different input sizes. The smaller input size is best on each task - especially on the landslide, where the damage extent is fine-grained - echoing the results on the transformer.
@@ VI-F2 Model Size
TABLE IV: Comparison of the transformer architecture with different sizes of feedforward layers (FF), number of layers (L), and resulting total number of parameters (P) on the disturbance mapping tasks. We report PR AUC for each disturbance mapping task. Best results are bolded. The 3.2 million parameter model is best across all tasks.
@@ VI-F3 Learning Rate
Finally, we ablate the learning rate. Table V compares the transformer model trained at different starting learning rates (with the same learning rate decay factor of 0.1 at epoch 25). The other modeling and training choices are the same as reported in Section III-C. The starting learning rate of 1e-4 exhibits the strongest results, which corresponds to the results we report in the main text.
@@
TABLE V: Comparison of the same transformer architectures trained with different starting learning rates (LRs) on the disturbance mapping tasks. We report PR AUC for each disturbance mapping task. Best results are bolded. LRs larger than 1e-4 resulted in little to no training and are omitted.
@@ VII Discussion
Although we are far from verifying this approach is globally applicable, the transformer was consistently the best performing method on three different disaster events in three different parts of the world. We also note that the optimal thresholds for the transformer all fall within $4$ to $7$ standard deviations (see Figure 18); in particular, $\tau\approx 5$ would produce $F_{1}\approx 0.6$ across all three sites. This indicates some inherent generality in the model, especially for operational purposes, where the threshold $\tau$ would have to be chosen without knowledge of ground truth. In practice, the choice of threshold will depend on the application: a lower threshold will capture more true positives (that is, disturbances) at the expense of possibly more false positives, while a higher threshold will ensure fewer false positives but risk missing some true positives.
@@ VIII Conclusion and Future Work
In this work, we trained deep self-supervised vision transformer on near-global OPERA RTC-S1 SAR data. Given a sequence of RTC-S1 baseline imagery, the transformer estimates a per-pixel distribution. By comparing this predicted distribution to an image outside of the baseline, we quantified disturbance via a transformer metric (i.e. the Mahalanobis distance) and showed that this transformer metric can be thresholded to delineate disturbance. We systematically analyzed this transformer-based metric to show it had strong agreement with externally validated data. While many previous works on SAR disturbance require supervision or are trained on a particular location and disaster event, our model was self-supervised without needing any fine-tuning for different environments or different categories of disaster.
There are two avenues of future work for this approach on disturbance. One avenue is to provide systematic verification that our vision transformer approach has global applicability for the RTC-S1 data. This analysis includes assessing additional categories of disturbances such as logging, urbanization, mining which may evolve beyond a single Sentinel-1 acquisition; analyzing the necessary baseline imagery that is required to delineate various categories of disturbances; and analyzing how disturbance delineations change over time and how such disturbances can be consistently tracked over a time-series. The goal of such systematic analysis is to adapt this transformer disturbance algorithm for near global disturbance monitoring through RTC-S1 inputs.
Another avenue of future work is to improve the model itself including increasing the model’s depth; training the model on more data; adding more nuanced spatiotemporal features as in [76, 43], including a masked pretraining step as in [34]; and increasing the window size so that the vision transformer can expand its spatial context. A larger input size, such as 224x224, could be desirable [41], given sufficient data and compute. The ultimate goal of this type of work is to generate a SAR foundation model - that is, a transformer-based model specifically designed for for SAR imagery that can be fine-tuned across various wavelengths (C-band, L-band for the coming NISAR mission) and across multiple downstream tasks including disturbance mapping, land cover use change, classification. Indeed, promising work has been done in this direction for multispectral data [78, 77, 79] and hybrid remote sensing data [43, 113]. We believe the community would benefit from a SAR-specific approach, especially using the new OPERA RTC-S1 product and the forthcoming NISAR products.

### v2

@@ VI-E Sensitivity to $\tau$
First, we evaluate the sensitivity of each method’s $F_{1}$ score to the threshold $\tau$ for the landslide event in Fig. 14. Natively, the deep methods’ threshold is in units of standard deviations. For the log ratio, the threshold is in units of decibels. However, to compare all the methods side-by-side, we plot the thresholds on the x-axis as a percentage of the maximum threshold (where no pixels would be labeled positive). As described previously, the transformer achieves the best $F_{1}$ score. However, this plot illuminates that the transformer is also less sensitive to the threshold $\tau$ compared to the RNN and log ratio at their respective peaks.
Furthermore, in Fig. 15, we show that the transformer achieves strong $F_{1}$ scores on all tasks at similar thresholds; a threshold of approximately $5$ standard deviations would result in $F_{1}$ scores around $0.6$ on all three tasks. While further validation is required, this is initial evidence that one could a priori select $\tau\approx 5$ and expect the transformer to produce high quality disturbance delineations across environments - even when ground truth is not available. This is a critical quality for a model to possess for operational deployment. Future work will focus on further corroborating this claim on more disturbance events.
@@ VI-F Ablation Studies
This subsection describes and empirically justifies some of the transformer and RNN hyperparameter choices made in this work. Throughout the ablation studies, we report PR AUC on each damage mapping task (landslide, fire, and flood) and bold the best result.
@@ VI-F1 Input and Patch Size
Table II compares the performance of the same transformer-based model with different input and patch sizes. The architecture and training procedure for each model in the table is the same as reported in Section III-C.
TABLE II: Effect of input (I) and patch sizes (P) on transformer performance
@@ VI-F2 Model Size
TABLE IV: Effect of width of feedforward layers (FF), number of layers (L), and total number of parameters (P) on transformer performance
@@ VI-F3 Learning Rate
Finally, we ablate the learning rate. Table V compares the transformer model trained at different starting learning rates (with the same learning rate decay factor of 0.1 at epoch 25). The other modeling and training choices are the same as reported in Section III-C. The starting learning rate of 1e-4 exhibits the strongest results, which corresponds to the results we report in the main text. Learning rates larger than 1e-4 resulted in little to no training and are omitted.
TABLE V: Effect of learning rate on transformer performance.
@@ VII Discussion
Although we are far from verifying this approach is globally applicable, the transformer was consistently the best performing method on three different disaster events in three different parts of the world. We also note that the optimal thresholds for the transformer all fall within $4$ to $7$ standard deviations (see Fig. 15); in particular, $\tau\approx 5$ would produce $F_{1}\approx 0.6$ across all three sites. This indicates some inherent generality in the model, especially for operational purposes, where the threshold $\tau$ would have to be chosen without knowledge of ground truth. In practice, the choice of threshold will depend on the application: a lower threshold will capture more true positives (that is, disturbances) at the expense of possibly more false positives, while a higher threshold will ensure fewer false positives but risk missing some true positives.
@@ VIII Conclusion and Future Work
In this work, we trained deep self-supervised vision transformer on near-global OPERA RTC-S1 SAR data. We systematically analyzed this transformer-based metric to show it had strong agreement with externally validated data. While many previous works on SAR disturbance require supervision or are trained on a particular location and disaster event, our model was self-supervised without needing any fine-tuning for different environments or different categories of disaster.
There are two main avenues of future work. One is to provide systematic verification that our vision transformer approach has global applicability for the RTC-S1 data. This analysis includes assessing additional categories of disturbances such as logging, urbanization, and mining; analyzing the necessary baseline imagery required to delineate disturbances; and analyzing how disturbance delineations change and can be tracked over time. The goal of such systematic analysis is to adapt this transformer disturbance algorithm for near-global disturbance monitoring through RTC-S1 inputs.
Another avenue of future work is to improve the model itself including increasing the model’s depth, training the model on more data, adding more nuanced spatiotemporal features [83, 45], a masked pretraining step [36], and increasing the window size so that the vision transformer can expand its spatial context. The ultimate goal of this type of work is to generate a SAR foundation model that can be fine-tuned across various wavelengths (such as C-band and L-band for the coming NISAR mission) and across multiple downstream tasks including disturbance mapping, land cover use change, classification. Indeed, promising work has been done in this direction for multispectral data [85, 84, 86] and hybrid remote sensing data [45, 120]. We believe the community would benefit from a SAR-specific approach, especially using the new OPERA RTC-S1 product and the forthcoming NISAR products.